Denise bought 50 postage stamps for which she paid $13.00. Some were 20-cent stamps and some were 40-cent stamps. How many of each did she buy?
Answer by checkley77(12569)

X+Y=50 OR X=50-Y
.20X+.40Y=13
.20(50-Y)+.40Y=13
10-.20Y+.40Y=13
.20Y=13-10
.20Y=3
Y=3/.20
Y=15 40 CENT STAMPS WERE BOUGHT.
50-15=35 20 CENT STAMPS WERE BOUGHT.
PROOF:
.20*35+.40*15=13
7+6=13
13=13